### v2

@@ Abstract
This paper considers the network slicing (NS) problem which attempts to map multiple customized virtual network requests to a common shared network infrastructure and allocate network resources to meet diverse service requirements. This paper proposes an efficient decomposition algorithm for globally solving the large-scale NP-hard NS problem. The proposed algorithm decomposes the hard NS problem into two relatively easy function placement (FP) and traffic routing (TR) subproblems and iteratively solves them enabling the information feedback between each other, which makes it particularly suitable to solve large-scale problems. Specifically, the FP subproblem is to place service functions into cloud nodes in the network, and solving it can return a function placement strategy based on which the TR subproblem is defined; and the TR subproblem is to find paths connecting two nodes hosting two adjacent functions in the network, and solving it can either verify that the solution of the FP subproblem is an optimal solution of the original problem, or return a valid inequality to the FP subproblem that cuts off the current infeasible solution. The proposed algorithm is guaranteed to find the global solution of the NS problem. By taking the special structure of the NS problem into consideration, we successfully develop two families of valid inequalities that render the proposed decomposition algorithm converge much more quickly and thus much more efficient. We demonstrate the effectiveness and efficiency of the proposed valid inequalities and algorithm via numerical experiments.
@@
Decomposition algorithm, Farkas’ Lemma, network slicing, resource allocation, valid inequality.
@@ I-A Literature Review
To quickly obtain a feasible solution of the NS problem, various heuristic and meta-heuristic algorithms have also been proposed in the literature. In particular, references [16]-[24] proposed two-stage heuristic algorithms that first decompose the link-based MBLP formulation of the NS problem into a function placement (FP) subproblem and a traffic routing (TR) subproblem and then solve each subproblem in a one-shot fashion. The FP and TR subproblems attempt to map service functions into cloud nodes and find paths connecting two nodes hosting two adjacent functions in the network, respectively. To solve the FP subproblem, the works [16]-[18] first solved the linear programming (LP) relaxation of the NS problem and then used a rounding strategy while the works [19]-[24] used some greedy heuristics (without solving any LP). Once a solution for the mapping of service functions and cloud nodes is obtained, the TR subproblem can be solved by using shortest path, $k$-shortest path, or multicommodity flow algorithms. Unfortunately, solving the FP subproblem in a heuristic manner without taking the TR subproblem into account can lead to infeasibility or low-quality solutions. The work [25] divided the services into several groups to obtain several “small” subproblems, and sequentially solved the subproblems using standard MBLP solvers. Similarly, as this approach failed to simultaneously take the function placement and traffic routing of all services into account, it is likely to lead to infeasibility or return low-quality solutions. The works [9]-[14] proposed column generation (CG) approaches [30] to solve the path-based MBLP formulation of the NS problem. Unfortunately, only a subset of the paths is considered for the traffic routing of flows in the CG approach, which generally leads to infeasibility or low-quality solutions. Meta-heuristic algorithms were also developed to obtain a feasible solution for the NS problem including particle swarm optimization [26], simulated annealing [27], and Tabu search [28, 29].
In summary, the existing approaches to solving the NS problem either suffer from low efficiency (as they need to call MBLP solvers to solve MBLP formulations with a large problem size) or return a low-quality solution (as they fail to take the information of the whole problem into consideration). The goal of this paper is to fill this research gap, i.e., develop an algorithm that finds an optimal solution for the large-scale NS problem while still enjoying high computational efficiency.
@@ I-B Our contributions
The main contribution of this paper is the proposed decomposition algorithm for solving large-scale NS problems, which decomposes the hard large-scale NS problem into two relatively easy FP and TR subproblems and iteratively solves them enabling the information feedback between each other. Two key features of the proposed algorithm are: (i) It is guaranteed to find an optimal solution of the NS problem if the number of iterations between the FP and TR subproblems is allowed to be sufficiently large. Even though the number of iterations between the subproblems is small (e.g., 5), the proposed algorithm can still return a much better solution than existing two-stage heuristic algorithms [16]-[24]. (ii) The two subproblems in the proposed decomposition algorithm are much easier to solve than the original problem, which renders it particularly suitable to solve large-scale NS problems, compared with existing algorithms based on LP relaxations [16, 17] and algorithms needed to call MBLP solvers [7].
@@
  • Different from existing two-stage heuristic algorithms [16]-[24] that solve the FP and TR subproblems in a one-shot fashion, our proposed decomposition algorithm solves the two subproblems in an iterative fashion enabling the information feedback between them. The information feedback between the two subproblems plays a central role in the decomposition algorithm, which is the first technical contribution of the paper. In particular, solving the FP subproblem can return a function placement strategy based on which the TR subproblem is defined; and solving the TR subproblem can either verify that the solution of the FP subproblem is an optimal solution of the original problem or return a valid inequality to the FP subproblem that cuts off the current infeasible solution by virtue of Farkas’ lemma.
  • We propose two families of valid inequalities that judiciously exploit and utilize the special structures of the NS problem into consideration such as the connectivity and limited link capacity structures of the network. The proposed inequalities, as the second technical contribution of this paper, can be directly added in the FP subproblem, thereby significantly reducing the gap between the FP problem and the NS problem. Consequently, the proposed decomposition algorithm with these valid inequalities converges much faster and thus is much more efficient than that without the corresponding inequalities.
Numerical results demonstrate that the proposed valid inequalities significantly reduce the gap between the FP and NS problems and accelerate the convergence of the decomposition algorithm; and the proposed decomposition algorithm significantly outperforms the existing SOTA algorithms [7, 16, 17] in terms of both solution efficiency and quality.
In our prior work [1], we presented a decomposition framework for solving the NS problem. This paper is a significant extension of [1] towards a much more efficient decomposition algorithm. In particular, we develop two families of valid inequalities to accelerate the convergence of the decomposition algorithm. Moreover, we provide detailed computational results showing the effectiveness of the developed valid inequalities in reducing the gap and improving the performance of the decomposition algorithm. The development of valid inequalities and related simulation results are completely new compared to our prior work.
The rest of the paper is organized as follows. Section II introduces the system model and mathematical formulation for the NS problem. Sections III and IV present the decomposition algorithm and the valid inequalities to speed up the algorithm, respectively. Section V reports the numerical results. Finally, Section VI draws the conclusion.
@@ II System Model and Problem Formulation
Let $\mathcal{G}=\{\mathcal{I},\mathcal{L}\}$ denote the substrate (directed) network, where =$\mathcal{I}=\{i\}$ and =$\mathcal{L}=\{(i,j)\}$ are the sets of nodes and links, respectively. Let $\mathcal{V}\subseteq\mathcal{I}$ be the set of cloud nodes that can process functions. As assumed in [17], processing one unit of data rate consumes one unit of (normalized) computational capacity, and the total data rate on cloud node $v\in\mathcal{V}$ must be upper bounded by its capacity $\mu_{v}$. Similarly, each link $(i,j)\in\mathcal{L}$ has a communication capacity $C_{ij}$. There is a set of services $\mathcal{K}=\{k\}$ that need to be supported by the network. Each service $k\in\mathcal{K}$ relates to a customized service function chain (SFC) consisting of $\ell_{k}$ service functions that have to be processed in sequence by the network: $f_{1}^{k}\rightarrow f_{2}^{k}\rightarrow\cdots\rightarrow f_{\ell_{k}}^{k}$ [31]. In order to minimize the coordination overhead, each function must be processed at exactly one cloud node [17, 21]. For service $k$, let $S^{k},D^{k}\notin\mathcal{V}$ denote the source and destination nodes, and let $\lambda^{k}_{0}$ and $\lambda^{k}_{s}$, $s\in\mathcal{F}^{k}:=\{1,2,\ldots,\ell_{k}\}$, denote the data rates before receiving any function and after receiving function $f^{k}_{s}$, respectively.
The NS problem is to determine the function placement, the routes, and the associated data rates on the corresponding routes of all services while satisfying the capacity constraints of all cloud nodes and links. Next, we shall briefly introduce the problem formulation; see [17] for more details.
@@
We use $(k,s)$, $1\leq s<\ell_{k}$, to denote the traffic flow which is routed between the two cloud nodes hosting the two adjacent functions $f_{s}^{k}$ and $f_{s+1}^{k}$ (with the date rate being $\lambda_{s}^{k}$). Similarly, $(k,0)$ denotes the traffic flow which is routed between the source $S^{k}$ and the cloud node hosting function $f_{1}^{k}$ (with the date rate being $\lambda_{0}^{k}$), and $(k,\ell_{k})$ denotes the traffic flow which is routed between the cloud node hosting function $f_{\ell_{k}}^{k}$ and the destination $D^{k}$ (with the date rate being $\lambda_{\ell_{k}}^{k}$). Let $r_{ij}^{k,s}$ be the fraction of data rate $\lambda_{s}^{k}$ of flow $(k,s)$ on link $(i,j)$. Then the link capacity constraints can be written as follows:
@@
$\min_{\begin{subarray}{c}\bm{x},\,\bm{y},\,\bm{r}\end{subarray}}~{}$ $\sum_{v\in\mathcal{V}}p_{v}y_{v}+\sum_{v\in\mathcal{V}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}}c^{k,s}_{v}x^{k,s}_{v}$ (NS)
@@
III An Efficient Decomposition Algorithm
@@ III-A Proposed Algorithm
ptFP subproblemptTR subproblemdeeply interwoven witheach other via constraint (5)
Figure 1: Problem (NS) is a combination of the FP and TR subproblems and the two subproblems are deeply interwoven with each other via constraint (5).
Before going into details, we give a high-level preview of the proposed decomposition algorithm. From Section II, problem (NS) is a combination of the FP subproblem and the TR subproblem and the two subproblems are deeply interwoven with each other via constraints (5), where the placement variable $\bm{x}$ and the traffic routing variable $\bm{r}$ are coupled; see Fig. 1 for an illustration. The FP subproblem addresses the placement of functions into cloud nodes in the network while the TR subproblem addresses the traffic routing of all pairs of two cloud nodes hosting two adjacent functions. The basic idea of the proposed decomposition algorithm is to decompose the “hard” problem (NS) into two relatively “easy” subproblems, which are the FP and TR subproblems, and iteratively solve them with a useful information feedback between the subproblems (to deal with the coupled constraints) until an optimal solution of problem (NS) is found. In particular, solving the FP subproblem can provide a solution $\bar{\bm{x}}$ to the TR subproblem based on which the TR subproblem is defined; and solving the TR subproblem can either verify that $\bar{\bm{x}}$ is an optimal solution of problem (NS), or return a valid inequality of the form $\bm{\pi}^{\top}\bm{x}\geq\pi_{0}$ for the FP subproblem that cuts off the current infeasible solution $\bar{\bm{x}}$. The information feedback between the FP and TR subproblems in the decomposition algorithm is illustrated in Fig. 2. In the following, we shall present the decomposition algorithm in detail.
ptFP subproblemptTR subproblem
@@ III-A1 FP subproblem
$\min_{\begin{subarray}{c}\bm{x},\,\bm{y}\end{subarray}}~{}$ $\sum_{v\in\mathcal{V}}p_{v}y_{v}+\sum_{v\in\mathcal{V}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}}c^{k,s}_{v}x^{k,s}_{v}$ (FP)
@@
Problem (FP) is an MBLP problem and is a relaxation of problem (NS) in which all traffic routing related variables (i.e., $\bm{r}$) and constraints (i.e., (4), (5), and (8)) are dropped. Note that the dimension of problem (FP) is significantly smaller than that of problem (NS). Therefore, solving problem (FP) by the SOTA MBLP solvers is significantly more efficient than solving problem (NS). To determine whether a solution $(\bar{\bm{x}},\bar{\bm{y}})$ of problem (FP) can define an optimal solution of problem (NS), i.e., whether there exists a point $\bar{\bm{r}}$ such that $(\bar{\bm{x}},\bar{\bm{y}},\bar{\bm{r}})$ is a feasible solution of problem (NS), we need to solve the TR subproblem.

### v3

@@ Abstract
This paper considers the network slicing (NS) problem which attempts to map multiple customized virtual network requests to a common shared network infrastructure and allocate network resources to meet diverse service requirements. This paper proposes an efficient customized Benders decomposition algorithm for globally solving the large-scale NP-hard NS problem. The proposed algorithm decomposes the hard NS problem into two relatively easy function placement (FP) and traffic routing (TR) subproblems and iteratively solves them enabling the information feedback between each other, which makes it particularly suitable to solve large-scale problems. Specifically, the FP subproblem is to place service functions into cloud nodes in the network, and solving it can return a function placement strategy based on which the TR subproblem is defined; and the TR subproblem is to find paths connecting two nodes hosting two adjacent functions in the network, and solving it can either verify that the solution of the FP subproblem is an optimal solution of the original problem, or return a valid inequality to the FP subproblem that cuts off the current infeasible solution. The proposed algorithm is guaranteed to find the globally optimal solution of the NS problem. By taking the special structure of the NS problem into consideration, we successfully develop two families of valid inequalities that render the proposed algorithm converge much more quickly and thus much more efficient. Numerical results demonstrate that the proposed valid inequalities effectively accelerate the convergence of the decomposition algorithm, and the proposed algorithm significantly outperforms the existing algorithms in terms of both solution efficiency and quality.
@@
Benders decomposition, Farkas’ Lemma, network slicing, resource allocation, valid inequality.
@@ I-A Literature Review
To quickly obtain a feasible solution of the NS problem, various heuristic and meta-heuristic algorithms have also been proposed in the literature. In particular, the works [16]-[24] proposed two-stage heuristic algorithms that first decompose the link-based MBLP formulation of the NS problem into a function placement (FP) subproblem and a traffic routing (TR) subproblem and then solve each subproblem in a one-shot fashion. The FP and TR subproblems attempt to map service functions into cloud nodes and find paths connecting two nodes hosting two adjacent functions in the network, respectively. To solve the FP subproblem, the work [16] first solved the linear programming (LP) relaxation of the NS problem and then used a one-shot rounding strategy; the work [17] used an LP dynamic rounding algorithm that sequentially rounds the value of a variable so that its value is consistent to other already rounded variables; the work [18] first solved a series of LP relaxation problems to find a near integral solution and then used a one-shot rounding strategy; and the works [19]-[24] used some greedy heuristics (without solving any LP). Once a solution for the mapping of service functions and cloud nodes is obtained, the TR subproblem can be solved by using shortest path, $k$-shortest path, or multicommodity flow algorithms. Unfortunately, solving the FP subproblem in a heuristic manner without taking the TR subproblem into account can lead to infeasibility or low-quality solutions. The work [25] divided the services into several groups to obtain several “small” subproblems, and sequentially solved the subproblems using standard MBLP solvers. Similarly, as this approach failed to simultaneously take the function placement and traffic routing of all services into account, it is likely to lead to infeasibility or return low-quality solutions. The works [9]-[14] proposed column generation (CG) approaches [30] to solve the path-based MBLP formulation of the NS problem. Unfortunately, only a subset of the paths is considered for the traffic routing of flows in the CG approach, which generally leads to infeasibility or low-quality solutions. Meta-heuristic algorithms were also developed to obtain a feasible solution for the NS problem including particle swarm optimization [26], simulated annealing [27], and Tabu search [28, 29].
In summary, the existing approaches to solving the NS problem either suffer from low efficiency (as they need to call MBLP solvers to solve MBLP formulations with a large problem size) or return a low-quality solution (as they fail to take the information of the whole problem into consideration). The main motivation of this paper is to develop an efficient algorithm that takes the information of the whole problem into consideration and finds a globally optimal solution for the large-scale NS problem. Integer programming and Benders decomposition techniques [30] play vital roles in the developed algorithm.
@@ I-B Our contributions
The main contribution of this paper is the proposed customized Benders decomposition (CBD) algorithm for solving large-scale NS problems, which decomposes the hard large-scale NS problem into two relatively easy FP and TR subproblems and iteratively solves them enabling the information feedback between each other. Two key features of the proposed algorithm are: (i) It is guaranteed to find an optimal solution of the NS problem if the number of iterations between the FP and TR subproblems is allowed to be sufficiently large. Even though the number of iterations between the subproblems is small (e.g., 5), the proposed algorithm can still return a much better solution than existing two-stage heuristic algorithms [16]-[24]. (ii) The two subproblems in the proposed decomposition algorithm are much easier to solve than the original problem, which renders it particularly suitable to solve large-scale NS problems, compared with existing algorithms based on LP relaxations [16, 17, 18] and algorithms needed to call MBLP solvers [7].
@@
  • Different from existing two-stage heuristic algorithms [16]-[24] that solve the FP and TR subproblems in a one-shot fashion, our proposed decomposition algorithm solves the two subproblems in an iterative fashion enabling the information feedback between them. The information feedback between the two subproblems plays a central role in the proposed algorithm, which is the first technical contribution of the paper. In particular, solving the FP subproblem can return a function placement strategy based on which the TR subproblem is defined; and solving the TR subproblem can either verify that the solution of the FP subproblem is an optimal solution of the original problem or return a valid inequality to the FP subproblem that cuts off the current infeasible solution by virtue of Farkas’ lemma.
  • We propose two families of valid inequalities that judiciously exploit and utilize the special structures of the NS problem into consideration such as the connectivity and limited link capacity structures of the network. The proposed inequalities, as the second technical contribution of this paper, can be directly added in the FP subproblem, thereby significantly reducing the gap between the FP problem and the NS problem. Consequently, the proposed algorithm with these valid inequalities converges much faster and thus is much more efficient than that without the corresponding inequalities.
Numerical results demonstrate that the proposed valid inequalities significantly reduce the gap between the FP and NS problems and accelerate the convergence of the decomposition algorithm; and the proposed decomposition algorithm significantly outperforms the existing SOTA algorithms [7, 16, 17, 18] in terms of both solution efficiency and quality.
In our prior work [1], we presented a CBD framework for solving the NS problem. This paper is a significant extension of [1] towards a much more efficient CBD algorithm. In particular, we develop two families of valid inequalities to accelerate the convergence of the CBD algorithm. Moreover, we provide detailed computational results showing the effectiveness of the developed valid inequalities in reducing the gap and improving the performance of the CBD algorithm. The development of valid inequalities and related simulation results are completely new compared to our prior work.
The rest of the paper is organized as follows. Section II introduces the system model and mathematical formulation for the NS problem. Sections III and IV present the CBD algorithm and the valid inequalities to speed up the algorithm, respectively. Section V reports the numerical results. Finally, Section VI draws the conclusion.
@@ II System Model and Problem Formulation
Let $\mathcal{G}=\{\mathcal{I},\mathcal{L}\}$ denote the substrate (directed) network, where =$\mathcal{I}=\{i\}$ and =$\mathcal{L}=\{(i,j)\}$ are the sets of nodes and links, respectively. Let $\mathcal{V}\subseteq\mathcal{I}$ be the set of cloud nodes that can process functions. As assumed in [18], processing one unit of data rate consumes one unit of (normalized) computational capacity, and the total data rate on cloud node $v\in\mathcal{V}$ must be upper bounded by its capacity $\mu_{v}$. Similarly, each link $(i,j)\in\mathcal{L}$ has a communication capacity $C_{ij}$. There is a set of services $\mathcal{K}=\{k\}$ that need to be supported by the network. Each service $k\in\mathcal{K}$ relates to a customized service function chain (SFC) consisting of $\ell_{k}$ service functions that have to be processed in sequence by the network: $f_{1}^{k}\rightarrow f_{2}^{k}\rightarrow\cdots\rightarrow f_{\ell_{k}}^{k}$ [31]. If a function is processed by multiple nodes, it may introduce the overhead of maintaining the network state consistency between different parts of the function [32, 33] (as they need to communicate with each other to exchange data or synchronize their states). In order to reduce the coordination overhead, we follow [18, 21, 32, 33] to assume that each function must be processed at exactly one cloud node. For service $k$, let $S^{k},D^{k}\notin\mathcal{V}$ denote the source and destination nodes, and let $\lambda^{k}_{0}$ and $\lambda^{k}_{s}$, $s\in\mathcal{F}^{k}:=\{1,2,\ldots,\ell_{k}\}$, denote the data rates before receiving any function and after receiving function $f^{k}_{s}$, respectively. In this paper, we assume that the service functions $\{f_{s}^{k}\}$ and their data rates $\{\lambda_{s}^{k}\}$ are known before mapping them into the substrate network, which is a common assumption in the literature [2, 11, 21].
The NS problem is to determine the function placement, the routes, and the associated data rates on the corresponding routes of all services while satisfying the capacity constraints of all cloud nodes and links. Next, we shall introduce the problem formulation.
@@
We use $(k,s)$, $1\leq s<\ell_{k}$, to denote the traffic flow which is routed between the two cloud nodes hosting the two adjacent functions $f_{s}^{k}$ and $f_{s+1}^{k}$ (with the data rate being $\lambda_{s}^{k}$). Similarly, $(k,0)$ denotes the traffic flow which is routed between the source $S^{k}$ and the cloud node hosting function $f_{1}^{k}$ (with the data rate being $\lambda_{0}^{k}$); $(k,\ell_{k})$ denotes the traffic flow which is routed between the cloud node hosting function $f_{\ell_{k}}^{k}$ and the destination $D^{k}$ (with the date rate being $\lambda_{\ell_{k}}^{k}$). Let $r_{ij}^{k,s}$ be the fraction of data rate $\lambda_{s}^{k}$ of flow $(k,s)$ on link $(i,j)$. Then the link capacity constraints can be written as follows:
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\operatorname*{minimize}_{\begin{subarray}{c}\bm{x},\,\bm{y},\,\bm{r}\end{subarray}}}}{}$ $\sum_{v\in\mathcal{V}}p_{v}y_{v}+\sum_{v\in\mathcal{V}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}}c^{k,s}_{v}x^{k,s}_{v}$ (NS)
@@
III An Efficient CBD Algorithm
Benders decomposition is an algorithmic framework that can globally solve large-scale mixed integer linear programming problems [30]. Its basic idea is to decompose the original problem into two simpler problems, called Benders master and subproblems, and iteratively solve the two subproblems until an optimal solution of the original problem is found. In this section, we shall develop an efficient CBD algorithm for solving the large-scale NS problem based on its special decomposition structure. We shall also provide some analysis results on the proposed algorithm.
@@ III-A Proposed Algorithm
Before going into details, we first give a high-level preview of the proposed decomposition algorithm for solving problem (NS). From Section II, we can see the decomposition structure of problem (NS). More specifically, problem (NS) is a combination of the FP subproblem and the TR subproblem and the two subproblems are deeply interwoven with each other via constraints in (5), where the placement variable $\bm{x}$ and the traffic routing variable $\bm{r}$ are coupled. The FP subproblem addresses the placement of functions into cloud nodes in the network while the TR subproblem addresses the traffic routing of all pairs of two cloud nodes hosting two adjacent functions. The basic idea of the proposed decomposition algorithm is to first exploit the decomposition structure of problem (NS) and decompose the “hard” problem (NS) into two relatively “easy” subproblems, and then to iteratively solve them with a useful information feedback between the subproblems (to deal with the coupled constraints) until an optimal solution of problem (NS) is found. The FP and TR subproblems correspond to the Benders master and subproblems, respectively. In particular, solving the FP subproblem can provide a solution $\bar{\bm{x}}$ to the TR subproblem based on which the TR subproblem is defined; and solving the TR subproblem can either verify that $\bar{\bm{x}}$ is an optimal solution of problem (NS), or return a valid inequality of the form $\bm{\pi}^{\top}\bm{x}\geq\pi_{0}$ (called Benders cut) for the FP subproblem that cuts off the current infeasible solution $\bar{\bm{x}}$. Here, $\bm{\pi}$ and $\pi_{0}$ are the coefficient vector of the variable $\bm{x}$ and the right-hand side of the valid inequality, respectively, which can be obtained by solving the TR subproblem. The information feedback between the FP and TR subproblems in the proposed algorithm is illustrated in Fig. 1. In the following, we shall present the algorithm in detail.
FP subproblemTR subproblem
@@ III-A1 FP subproblem
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\operatorname*{minimize}_{\begin{subarray}{c}\bm{x},\,\bm{y}\end{subarray}}}}{}$ $\sum_{v\in\mathcal{V}}p_{v}y_{v}+\sum_{v\in\mathcal{V}}\sum_{k\in\mathcal{K}}\sum_{s\in\mathcal{F}^{k}}c^{k,s}_{v}x^{k,s}_{v}$ (FP)
@@
Problem (FP) is an MBLP problem and is a relaxation of problem (NS) in which all traffic routing related variables (i.e., $\bm{r}$) and constraints (i.e., (4), (5), and (8)) are dropped. Note that the dimension of problem (FP) is significantly smaller than that of problem (NS); see next subsection for a detailed discussion. Therefore, solving problem (FP) by the SOTA MBLP solvers is significantly more efficient than solving problem (NS). To determine whether a solution $(\bar{\bm{x}},\bar{\bm{y}})$ of problem (FP) can define an optimal solution of problem (NS), i.e., whether there exists a point $\bar{\bm{r}}$ such that $(\bar{\bm{x}},\bar{\bm{y}},\bar{\bm{r}})$ is a feasible solution of problem (NS), we need to solve the TR subproblem.